You are given the following information about a closed economy with no government:

Consumption = 445 + 0.75Y

Investment = 250

Use the above information to answer the questions that follow:

Q.4.1 Calculate the value of autonomous spending


1
Expert's answer
2021-10-25T09:50:32-0400

Let's find equilibrium income first:

AS"=" AD

Y"=" C"+" I

"Y=445+0.75Y+I"

"Y-0.75Y=445+250"

"0.25Y=695"

"Y=2780"

Consumption function is "C=445+0.75Y" where Y is the income in the economy. So at national income 2780, consumption spending is:

"C=445+0.75Y"

"=445+0.75(2780)"

"=445+2085"

"=2530"